OpenStudy (anonymous):

What is the value of i 20+1? 1 –1 –i i

OpenStudy (anonymous):

I got 2 but thats not an option?

OpenStudy (solomonzelman):

maybe you meant \(\large\color{black}{ \displaystyle i^{20+1} }\) ?

OpenStudy (anonymous):

yes! sorry

OpenStudy (solomonzelman):

yes, it is alright. next when you want to write that, just say i^(20+1) where ^ indicates an exponent, and (...) tell you what your exponent is.

OpenStudy (solomonzelman):

Anyway,

OpenStudy (solomonzelman):

Can you tell me what does \(\large\color{black}{ \displaystyle i^4 }\) equal to?

OpenStudy (anonymous):

oh duh, answer is i

OpenStudy (solomonzelman):

yes, that is 1.

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \displaystyle i^{20+1}=i^{20} \times i^1=i^{4\times 5}\times i=\left(i^4\right)^5\times i = 1^5\times i=? }\)

OpenStudy (solomonzelman):

well, you probably know anyway that \(\large\color{black}{ \displaystyle i^{4n}=1 }\) \(\large\color{black}{ \displaystyle i^{4n+1}=i }\) \(\large\color{black}{ \displaystyle i^{4n+2}=-1 }\) \(\large\color{black}{ \displaystyle i^{4n+3}=-i }\) and again restarting the cycle, \(\large\color{black}{ \displaystyle i^{4n+4}=1 }\) and so on... (this is true for any whole number powers of i)

OpenStudy (solomonzelman):

(saying, for any whole number n)

OpenStudy (solomonzelman):

If you got any questions, please ask...